# HQ-VAE: Hierarchical Discrete Representation Learning with Variational Bayes

### v1

@@ Appendix D Experimental details
CelebA-HQ. CelebA-HQ (Karras et al., 2018) contains 30,000 high-resolution face images that are selected from the CelebA dataset by following Karras et al. (2018). We use the default training/validation/test split. We preprocess the images by cropping and resizing them to 256$\times$256.
@@
ImageNet. ImageNet (Deng et al., 2009) contains 1000 classes of natural images in RGB scales. We use the default training/validation/test split. We crop and resize the images to 256$\times$256.
@@
D.1 SQ-VAE-2 vs VQ-VAE-2
D.1.1 Comparison on CIFAR10 and CelebA-HQ
@@
We construct the architecture as depicted in Figures 1 and 7. To build the top-down paths, we use two injected top-down layers (i.e., $R=2$) with $w_{1}=h_{1}=8$ and $w_{2}=h_{2}=16$ for CIFAR10, and three layers (i.e., $R=3$) with $w_{1}=h_{1}=8$, $w_{2}=h_{2}=16$ and $w_{3}=h_{3}=32$ for CelebA-HQ. For the bottom-up paths, we repeatedly stack two Resblocks and an average pooling layer once and four times, respectively, for CIFAR10 and CelebA-HQ. We set the learning rate to 0.001 and train all the models for a maximum of 100 epochs with a mini-batch size of 32.
D.1.2 Comparison on large-scale datasets
@@
D.2 RSQ-VAE vs RQ-VAE
D.2.1 Comparison on CIFAR10 and CelebA-HQ
We construct the architecture as depicted in Figures 1 and 7 without injected top-down layers, i.e., $R=1$. We set the resolution of $\bm{H}_{\bm{\phi}}(\bm{x})$ to $w=h=8$. For the bottom-up paths, we stack two Resblocks on the average pooling layer once for CIFAR10 and four times for CelebA-HQ. We set the learning rate to 0.001 and train all the models for a maximum of 100 epochs with a mini-batch size of 32.
D.2.2 Improvement in perceptual quality
This experiment use the same network architecture as in Lee et al. (2022a). We set the learning rate to 0.001 and train an RSQ-VAE model for a maximum of 300 epochs with a mini-batch size of 128 (4 GPUs, 32 samples for each GPU) on FFHQ. We use our modified LPIPS loss (see Appendix D.5) in the training. For the evaluation, we compute rFID scores with the code provided in their repository [^3] https://github.com/kakaobrain/rq-vae-transformer on the validation set (10,000 samples). Moreover, we use the pre-trained RQ-VAE model offered in the same repository for evaluating RQ-VAE.
We will show examples of reconstructed images in Appendix D.5 after we explain our modified LPIPS loss.
(a) $L=4$
(a) $L=4$
(b) $L=8$
@@ D.2.2 Improvement in perceptual quality
D.2.3 Validation on an audio dataset
@@
D.3 Empirical study of top-down layers
@@
Figure 11: Architecture of the hybrid model in Appendix D.3.
@@
D.4 Applications of HQ-VAEs to generative tasks
@@
D.5 Perceptual loss for images
@@ D.6 Progressive coding
$\mathcal{J}_{\text{SQ-VAE-2}}^{\text{prog}}=\sum_{l=1}^{L}\frac{D}{2}\log\sigma_{l}^{2}+\mathbb{E}_{\mathcal{Q}(\bm{Z}_{1:L},\tilde{\bm{Z}}_{1:L}|\bm{x})}\left[\frac{\|\bm{x}-f_{\bm{\theta}}(\bm{Z}_{1:l})+Dv_{l}\|_{2}^{2}}{2\sigma_{l}^{2}}+\frac{\|\tilde{\bm{Z}}_{l}-\bm{Z}_{l}\|_{F}^{2}}{2s_{l}^{2}}-H(\hat{P}_{s_{l}^{2}}(\bm{Z}_{l}|\tilde{\bm{Z}}_{l}))\right]$ (28)
@@
This concept can be also applied to the hybrid model derived in Appendix C by considering additional generative processes $p_{\bm{\theta}}^{r}(\tilde{\bm{x}}_{r})=\mathcal{N}(\tilde{\bm{x}}_{r};f_{\bm{\theta}}(\bm{Z}_{1:L_{r}}),\sigma_{r}^{2}\bm{I})$. The ELBO objective is as follows:
@@
In Section D.3, we simply set $v_{l}=0$ in the above objective when this technique is activated.
(a) Source.
(b) RQ-VAE with 4 layers.
(c) RQ-VAE with 8 layers.
(d) RSQ-VAE with 4 layers.
@@
(a) Source
(b) RSQ-VAE trained with a normal LPIPS loss (rFID$=10.07$)

### v2

@@ Appendix F Experimental details
CelebA-HQ. CelebA-HQ (Karras et al., 2018) contains 30,000 high-resolution face images that are selected from the CelebA dataset by following Karras et al. (2018). We use the default training/validation/test split (24,183/2,993/2,824 samples). We preprocess the images by cropping and resizing them to 256$\times$256.
@@
ImageNet. ImageNet (Deng et al., 2009) contains 1000 classes of natural images in RGB scales. We use the default training/validation/test split (1,281,167/50,000/100,000 samples). We crop and resize the images to 256$\times$256.
@@
F.1 SQ-VAE-2 vs VQ-VAE-2
F.1.1 Comparison on CIFAR10 and CelebA-HQ
@@
(d) SQ-VAE-2 and VQ-VAE-2 on CIFAR10
(e) RSQ-VAE and RQ-VAE on CIFAR10
Figure 8: Impact of batch size on reconstruction of images in CIFAR10. (a) We set the codebook capacity for the discrete space to $(d_{l},K_{l})=(4^{2},512),(8^{2},512)$ with $L=2$. (b) We set the codebook capacity for the discrete space to $(d_{l},K_{l})=(8^{2},32),(8^{2},32)$ with $L=2$. For (a) and (b), we use the same architectures as in Sections 5.1 and 5.2, respectively. The plots suggest that batch size should be tuned to achieve better reconstruction performance for all the models.
We construct the architecture as depicted in Figures 1 and 7. To build the top-down paths, we use two injected top-down layers (i.e., $R=2$) with $w_{1}=h_{1}=8$ and $w_{2}=h_{2}=16$ for CIFAR10, and three layers (i.e., $R=3$) with $w_{1}=h_{1}=8$, $w_{2}=h_{2}=16$ and $w_{3}=h_{3}=32$ for CelebA-HQ. For the bottom-up paths, we repeatedly stack two Resblocks and an average pooling layer once and four times, respectively, for CIFAR10 and CelebA-HQ. We set the learning rate to 0.001 and train all the models for a maximum of 100 epochs with a mini-batch size of 32. The sensitivity of SQ-VAE-2 in terms of batch size is investigated in Figure 8.
F.1.2 Comparison on large-scale datasets
@@
F.2 RSQ-VAE vs RQ-VAE
F.2.1 Comparison on CIFAR10 and CelebA-HQ
We construct the architecture as depicted in Figures 1 and 7 without injected top-down layers, i.e., $R=1$. We set the resolution of $\bm{H}_{\bm{\phi}}(\bm{x})$ to $w=h=8$. For the bottom-up paths, we stack two Resblocks on the average pooling layer once for CIFAR10 and four times for CelebA-HQ. We set the learning rate to 0.001 and train all the models for a maximum of 100 epochs with a mini-batch size of 32. The sensitivity of RSQ-VAE in terms of batch size is investigated in Figure 8.
F.2.2 Improvement in perceptual quality
This experiment use the same network architecture as in Lee et al. (2022a). We set the learning rate to 0.001 and train an RSQ-VAE model for a maximum of 300 epochs with a mini-batch size of 128 (4 GPUs, 32 samples for each GPU) on FFHQ. We use our modified LPIPS loss (see Appendix F.5) in the training. For the evaluation, we compute rFID scores with the code provided in their repository [^3] https://github.com/kakaobrain/rq-vae-transformer on the validation set (10,000 samples). Moreover, we use the pre-trained RQ-VAE model offered in the same repository for evaluating RQ-VAE.
We will show examples of reconstructed images in Appendix F.5 after we explain our modified LPIPS loss.
(c) $L=4$
(d) $L=8$
@@ F.2.2 Improvement in perceptual quality
F.2.3 Validation on an audio dataset
@@
F.3 Empirical study of top-down layers
@@
Figure 12: Architecture of the hybrid model in Appendix F.3.
@@
F.4 Applications of HQ-VAEs to generative tasks
@@
F.5 Perceptual loss for images
@@ F.6 Progressive coding
$\mathcal{J}_{\text{SQ-VAE-2}}^{\text{prog}}=\sum_{l=1}^{L}\frac{D}{2}\log\sigma_{l}^{2}+\mathbb{E}_{\mathcal{Q}(\bm{Z}_{1:L},\tilde{\bm{Z}}_{1:L}|\bm{x})}\left[\frac{\|\bm{x}-f_{\bm{\theta}}(\bm{Z}_{1:l})+Dv_{l}\|_{2}^{2}}{2\sigma_{l}^{2}}+\frac{\|\tilde{\bm{Z}}_{l}-\bm{Z}_{l}\|_{F}^{2}}{2s_{l}^{2}}-H(\hat{P}_{s_{l}^{2}}(\bm{Z}_{l}|\tilde{\bm{Z}}_{l}))\right]{\color[rgb]{0,0,0}{{.}}}$ (28)
@@
This concept can be also applied to the hybrid model derived in Appendix D by considering additional generative processes $p_{\bm{\theta}}^{r}(\tilde{\bm{x}}_{r})=\mathcal{N}(\tilde{\bm{x}}_{r};f_{\bm{\theta}}(\bm{Z}_{1:L_{r}}),\sigma_{r}^{2}\bm{I})$. The ELBO objective is as follows:
@@
In Section F.3, we simply set $v_{l}=0$ in the above objective when this technique is activated.